OpenStudy (supercait10):

What makes the statement true? 8^-2 A. 64 B. 16 C. -1/64 D. 1/64?

OpenStudy (alexandervonhumboldt2):

a^(-b)=1/a^b use this rule

OpenStudy (alexandervonhumboldt2):

\[a^{-b}=\frac{ 1 }{ a^b }\]

OpenStudy (alexandervonhumboldt2):

\[8^{-2}=\frac{ 1 }{ 8^2 }=???\]